# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 7 Period sheaves
$\mathbb{B}_{\mathrm{crys}}|_{X_{\beta}}\longrightarrow{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet},\qquad\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\longrightarrow{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet}$
@@
The proof is local. We may assume that $\mathcal{X}_{\beta}=\mathrm{Spf}(R_{\beta})\subset\mathfrak{X}_{\beta}=\mathrm{Spf}(\mathfrak{R}_{\beta})$ admits a chart as in §4.1. As explained in [AI, §3.1.2] such a choice of chart defines an integral perfectoid extension $\widehat{\overline{R}}^{0}_{\beta}$ of $R^{0}_{\beta}$ and hence $\widehat{\overline{R}}_{\beta}$ of $R_{\beta}$ and a perfectoid open $\mathcal{U}_{\beta}$ of $X_{\mathrm{pket}}$. Fix compatible $p$-power roots $\underline{\pi}$, $\underline{X}_{1}$, $\cdots$, $\underline{X}_{a}$, $\underline{Y}_{1}$, $\cdots$, $\underline{Y}_{b}$ in $\widehat{\overline{R}}^{\flat}_{\beta}$ of $\pi$, $X_{1}$, $\cdots$, $X_{a}$, $Y_{1}$, $\cdots$, $Y_{b}$ in $\widehat{\overline{R}}_{\beta}$ such that $\underline{X}_{1}\cdots\underline{X}_{a}=\underline{\pi}$. We denote by [_][\_] their Teichmüller lifts in $\mathbf{W}\bigl(\widehat{\overline{R}}^{\flat}_{\beta})$. Arguing as in [AI, §3.4], especially Remark 3.26, one gets that $\mathbb{A}_{\mathrm{crys}}(\mathcal{U}_{\beta})\cong A_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ and
@@
arguing as in [AI, §3.4], Remak 3.26, for the second. From this also the second claim follows. ∎
@@
We may replace $\mathfrak{E}_{\beta}$ with $\mathcal{O}_{\mathfrak{X}_{\beta}}$ as $\mathfrak{E}$ is locally free on the logarithmic crystalline site of $\mathcal{X}_{k}$. One then argues as in the proof of [TT, Lemma 5.3] reducing the claims to caims in group cohomology, namely Theorem 6.12 of loc. cit. Its analogue in our context is [AI, Thm. 3.39] for the computation of $\mathrm{R}^{i}(w\circ\nu_{\mathrm{geo}})_{\ast}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)$ for $i\geq 0$ and [AI, Prop. 3.40] for the computation of $(w\circ\nu_{\mathrm{ar}})_{\ast}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)$.
@@ 8 Crystalline associated sheaves
We let $\mathcal{L}$ be a $\mathbb{Z}_{p}$-local system on $X_{\mathrm{ket}}$ with unipotent geometric monodromy along $D$ ([DLLZa, Def. 6.3.1 & Def. 6.3.7]). We write $\widehat{\mathcal{L}}$ for the associated sheaf of $\widehat{\mathbb{Z}}_{p}$-modules in $X_{\mathrm{pket}}$ as in [DLLZa, Def. 6.3.2]. Let $\mathcal{E}$ be a locally free $\mathcal{O}_{X}$-module, endowed with an integrable connection $\nabla$ and a descending filtration $\mathrm{Fil}^{\bullet}\mathcal{E}$ by locally free $\mathcal{O}_{X}$-modules such that Griffiths’ transversality holds and the graded pieces are locally free $\mathcal{O}_{X}$-modules as well.
@@
As $\mathcal{L}$ is assumed to have unipotent geometric monodromy along $D$, it follows from [DLLZb, Thm. 3.2.12] that $(\mathcal{E},\nabla)$ has nilpotent residues along $D$ in the sense of Definition 1.
@@
(1) Due to our assumption that $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ have unipotent geometric monodromy along $D$, the statement on being de Rham associated follows from [DLLZb, Thm. 3.2.12]. The fact that they are also crystalline associated follows easily from the definition.
(2) The pull-backs are de Rham associated due to [DLLZb, Thm. 3.2.7(4)]. The statement then follows from Lemma 8.2 as $f$ defines a morphism of sheaves $f^{-1}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)\to\mathcal{O}_{\mathfrak{Y}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ on $Y_{\mathrm{pket}}$ compatible with connections. ∎
@@
2) $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ is the Frobenius log-crystal relative to $(X,\alpha_{X},\overline{\mathcal{X}},\alpha_{\overline{\mathcal{X}}})$ extending the relative de Rham cohomology of $A\to U$ and the relative crystalline cohomology of the mod $p$ special fiber of $\mathcal{A}\to\mathcal{U}=\mathcal{X}\backslash\mathcal{D}$ relative to $\mathcal{U}$, constructed in [MP, §3.2].
@@
By [FC, Thm. VI.1.1] we can extend $A\to U$ to a proper and log smooth morphism $\overline{A}\to X$. Then $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla\bigr)$ is identified with the first logarithmic de Rham cohomology sheaf of $\overline{A}$ relative to $X$ as they both coincide with the canonical extension of the first logarithmic de Rham cohomology sheaf of $A$ relative to $U$. The fact that $\mathcal{L}$ and $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla\bigr)$ are de Rham associated follows then from [DLLZb, Theorem 3.2.7(5)]. Thanks to [TT, Theorem 5.5] they are also crystalline associated when we restrict to the rigid analytic fiber $\mathcal{U}_{K}$ associated to $\mathcal{U}$. In this case, using the compatibility with the comparison isomorphism for $p$-divisible groups and Dieudonné theory as in [SW, §3.5], the isomorphism $\xi_{\beta,\mathrm{crys}}$ over the pro-Kummer étale site of $\mathcal{U}_{K}\cap X_{\beta}$ is induced by a map
@@
Then evaluating $\xi_{\beta,\mathrm{crys}}^{\prime}$ at the affinoid perfectoid defined by $R_{\beta}\{f^{-1}\}\subset\overline{R}_{\beta}\{f^{-1}\}$ we get a $2g\times 2g$ matrix $M$ with values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)$. The claim follows if we show that $M$ takes values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ as $\xi_{\beta,\mathrm{crys}}^{\prime}$ is then defined over the pro-Kummer étale site of $X_{\beta}$. As $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ are $p$-adically complete and separated, it suffices to prove that the reduction $M_{n}$ of $M$ modulo $p^{n}$ takes values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p^{n})$ for every $n$. It follows from the discussion in [MP, Prop. A.2.2 & §A.2.6] that $\xi_{\beta,\mathrm{crys}}^{\prime}$ extends to the completion of $R_{\beta}$ at every maximal ideal ; in loc. cit. the author works over a local field but one checks that the proof of Prop. A.2.2 uses the constructions of §A.1 that hold over a complete local normal domain.
@@
Since $p$ is a regular element in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ by [AI, Cor. 3.27], arguing for each entry of $M$ and proceeding by induction on $n$, one reduces to prove the following : let $s$ be an element of $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)/(p)$ that belongs to $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\otimes_{R_{\beta}}\widehat{R}_{\beta,x}\bigr)/(p)$ for the completion $\widehat{R}_{\beta,x}$ of the localization $R_{\beta,x}$ at each maximal ideal of $R_{\beta}$, then $s\in\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p)$. Using the explicit description of $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p)$ provided in [AI, Lemma 3.25], at the end of page 231, as a free $\overline{R}_{\beta}/\pi\overline{R}_{\beta}$-module for some pseudouniformizer $\pi\in\mathcal{O}_{\mathbb{C}_{p}}$ (and similarly for $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}[f^{-1}]\bigr)/(p)$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\otimes_{R_{\beta}}\widehat{R}_{\beta,x}\bigr)/(p)$ as free $\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$-module and respectively as $(\overline{R}_{\beta}/\pi\overline{R}_{\beta})\otimes_{R_{\beta}}\widehat{R}_{\beta,x}$-module with the same basis), one reduces to the following claim. Let $s\in\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$ be such that it lies in $(\overline{R}_{\beta}/\pi\overline{R}_{\beta})\otimes_{R_{\beta}}\widehat{R}_{\beta,x}$ for every maximal ideal $x$ of $R_{\beta}$, then $s\in\overline{R}_{\beta}/\pi\overline{R}_{\beta}$. This is clear since the map $\overline{R}_{\beta}/\pi\overline{R}_{\beta}\to\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$ is injective by the regularity result proven above and using that $R_{\beta,x}\to\widehat{R}_{\beta,x}$ is faithfully flat. ∎
@@
Analogously define $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r)$ to be the object in the bounded derived category of sheaves on $X$ given by the total complex associated to the morphism of complexes $\iota^{c\text{-}\infty}\circ(1-\frac{\varphi}{p^{r}})\oplus\iota^{c\text{-}\infty}$ :
$\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\right)\to\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\right)\oplus\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathrm{Gr}^{r}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}\right).$
@@
Let $\mathrm{H}^{i}\bigl(\mathbf{Syn}(\mathcal{L},r)\bigr)$, resp. $\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r)\bigr)$ to be the $i$-th homology groups of the complexes $\mathbf{Syn}(\mathcal{L},r)$ and $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r)$ respectively.
@@
We have morphisms of complexes $\mathbf{Syn}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}(\mathcal{L},r)$ and $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r)$.
@@
It follows from Theorem 1 and Definition 9 and by [DLLZb, Cor. 2.4.2] that have a quasi isomorphism
@@
compatible with Frobenius. The same holds with support conditions, see [LLZ, Cor. 2.5.13] for the de Rham case and Theorem 1 in the crystalline case, noticing that the isomorphisms $\xi_{\mathrm{dR}}$ and $\xi_{\beta,\mathrm{crys}}$ of Definition 9 induce analogous and compatible isomorphisms $\xi_{\mathrm{dR}}^{c\text{-}\infty}$ and $\xi_{\beta,\mathrm{crys}}^{c\text{-}\infty}$ replacing $\mathcal{O}\mathbb{B}_{\mathrm{dR}}$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ with $\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}$. It follows that we have morphisms of complexes

### v2

@@ 7 Period sheaves
$\mathbb{B}_{\mathrm{crys}}|_{X_{\beta}}\longrightarrow{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet},\qquad\mathbb{B}_{\mathrm{crys}}^{c-\infty}\longrightarrow{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet}$
@@
The proof is local. We may assume that $\mathcal{X}_{\beta}=\mathrm{Spf}(R_{\beta})\subset\mathfrak{X}_{\beta}=\mathrm{Spf}(\mathfrak{R}_{\beta})$ admits a chart as in §4.1. As explained in [AI12, §3.1.2] such a choice of chart defines an integral perfectoid extension $\widehat{\overline{R}}^{0}_{\beta}$ of $R^{0}_{\beta}$ and hence $\widehat{\overline{R}}_{\beta}$ of $R_{\beta}$ and a perfectoid open $\mathcal{U}_{\beta}$ of $X_{\mathrm{pket}}$. Fix compatible $p$-power roots $\underline{\pi}$, $\underline{X}_{1}$, $\cdots$, $\underline{X}_{a}$, $\underline{Y}_{1}$, $\cdots$, $\underline{Y}_{b}$ in $\widehat{\overline{R}}^{\flat}_{\beta}$ of $\pi$, $X_{1}$, $\cdots$, $X_{a}$, $Y_{1}$, $\cdots$, $Y_{b}$ in $\widehat{\overline{R}}_{\beta}$ such that $\underline{X}_{1}\cdots\underline{X}_{a}=\underline{\pi}$. We denote by [_][\_] their Teichmüller lifts in $\mathbf{W}\bigl(\widehat{\overline{R}}^{\flat}_{\beta})$. Arguing as in [AI12, §3.4], especially Remark 3.26, one gets that $\mathbb{A}_{\mathrm{crys}}(\mathcal{U}_{\beta})\cong A_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ and
@@
arguing as in [AI12, §3.4], Remak 3.26, for the second. From this also the second claim follows. ∎
@@
We may replace $\mathfrak{E}_{\beta}$ with $\mathcal{O}_{\mathfrak{X}_{\beta}}$ as $\mathfrak{E}$ is locally free on the logarithmic crystalline site of $\mathcal{X}_{k}$. One then argues as in the proof of [TT19, Lemma 5.3] reducing the claims to caims in group cohomology, namely Theorem 6.12 of loc. cit. Its analogue in our context is [AI12, Thm. 3.39] for the computation of $\mathrm{R}^{i}(w\circ\nu_{\mathrm{geo}})_{\ast}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)$ for $i\geq 0$ and [AI12, Prop. 3.40] for the computation of $(w\circ\nu_{\mathrm{ar}})_{\ast}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)$.
@@ 8 Crystalline associated sheaves
We let $\mathcal{L}$ be a $\mathbb{Z}_{p}$-local system on $X_{\mathrm{ket}}$ with unipotent geometric monodromy along $D$ ([DLL+02, Def. 6.3.1 & Def. 6.3.7]). We write $\widehat{\mathcal{L}}$ for the associated sheaf of $\widehat{\mathbb{Z}}_{p}$-modules in $X_{\mathrm{pket}}$ as in [DLL+02, Def. 6.3.2]. Let $\mathcal{E}$ be a locally free $\mathcal{O}_{X}$-module, endowed with an integrable connection $\nabla$ and a descending filtration $\mathrm{Fil}^{\bullet}\mathcal{E}$ by locally free $\mathcal{O}_{X}$-modules such that Griffiths’ transversality holds and the graded pieces are locally free $\mathcal{O}_{X}$-modules as well.
@@
As $\mathcal{L}$ is assumed to have unipotent geometric monodromy along $D$, it follows from [DLL+23, Thm. 3.2.12] that $(\mathcal{E},\nabla)$ has nilpotent residues along $D$ in the sense of Definition 1.
@@
(1) Due to our assumption that $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ have unipotent geometric monodromy along $D$, the statement on being de Rham associated follows from [DLL+23, Thm. 3.2.12]. The fact that they are also crystalline associated follows easily from the definition.
(2) The pull-backs are de Rham associated due to [DLL+23, Thm. 3.2.7(4)]. The statement then follows from Lemma 8.2 as $f$ defines a morphism of sheaves $f^{-1}\left(\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\right)\to\mathcal{O}_{\mathfrak{Y}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ on $Y_{\mathrm{pket}}$ compatible with connections. ∎
@@
2) $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ is the Frobenius log-crystal relative to $(X,\alpha_{X},\overline{\mathcal{X}},\alpha_{\overline{\mathcal{X}}})$ extending the relative de Rham cohomology of $A\to U$ and the relative crystalline cohomology of the mod $p$ special fiber of $\mathcal{A}\to\mathcal{U}=\mathcal{X}\backslash\mathcal{D}$ relative to $\mathcal{U}$, constructed in [MAD19, §3.2].
@@
By [FC90, Thm. VI.1.1] we can extend $A\to U$ to a proper and log smooth morphism $\overline{A}\to X$. Then $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla\bigr)$ is identified with the first logarithmic de Rham cohomology sheaf of $\overline{A}$ relative to $X$ as they both coincide with the canonical extension of the first logarithmic de Rham cohomology sheaf of $A$ relative to $U$. The fact that $\mathcal{L}$ and $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla\bigr)$ are de Rham associated follows then from [DLL+23, Theorem 3.2.7(5)]. Thanks to [TT19, Theorem 5.5] they are also crystalline associated when we restrict to the rigid analytic fiber $\mathcal{U}_{K}$ associated to $\mathcal{U}$. In this case, using the compatibility with the comparison isomorphism for $p$-divisible groups and Dieudonné theory as in [SW13, §3.5], the isomorphism $\xi_{\beta,\mathrm{crys}}$ over the pro-Kummer étale site of $\mathcal{U}_{K}\cap X_{\beta}$ is induced by a map
@@
Then evaluating $\xi_{\beta,\mathrm{crys}}^{\prime}$ at the affinoid perfectoid defined by $R_{\beta}\{f^{-1}\}\subset\overline{R}_{\beta}\{f^{-1}\}$ we get a $2g\times 2g$ matrix $M$ with values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)$. The claim follows if we show that $M$ takes values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ as $\xi_{\beta,\mathrm{crys}}^{\prime}$ is then defined over the pro-Kummer étale site of $X_{\beta}$. As $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ are $p$-adically complete and separated, it suffices to prove that the reduction $M_{n}$ of $M$ modulo $p^{n}$ takes values in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p^{n})$ for every $n$. It follows from the discussion in [MAD19, Prop. A.2.2 & §A.2.6] that $\xi_{\beta,\mathrm{crys}}^{\prime}$ extends to the completion of $R_{\beta}$ at every maximal ideal ; in loc. cit. the author works over a local field but one checks that the proof of Prop. A.2.2 uses the constructions of §A.1 that hold over a complete local normal domain.
@@
Since $p$ is a regular element in $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$ by [AI12, Cor. 3.27], arguing for each entry of $M$ and proceeding by induction on $n$, one reduces to prove the following : let $s$ be an element of $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\overline{R}_{\beta}\{f^{-1}\}\bigr)/(p)$ that belongs to $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\otimes_{R_{\beta}}\widehat{R}_{\beta,x}\bigr)/(p)$ for the completion $\widehat{R}_{\beta,x}$ of the localization $R_{\beta,x}$ at each maximal ideal of $R_{\beta}$, then $s\in\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p)$. Using the explicit description of $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)/(p)$ provided in [AI12, Lemma 3.25], at the end of page 231, as a free $\overline{R}_{\beta}/\pi\overline{R}_{\beta}$-module for some pseudouniformizer $\pi\in\mathcal{O}_{\mathbb{C}_{p}}$ (and similarly for $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}[f^{-1}]\bigr)/(p)$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\otimes_{R_{\beta}}\widehat{R}_{\beta,x}\bigr)/(p)$ as free $\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$-module and respectively as $(\overline{R}_{\beta}/\pi\overline{R}_{\beta})\otimes_{R_{\beta}}\widehat{R}_{\beta,x}$-module with the same basis), one reduces to the following claim. Let $s\in\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$ be such that it lies in $(\overline{R}_{\beta}/\pi\overline{R}_{\beta})\otimes_{R_{\beta}}\widehat{R}_{\beta,x}$ for every maximal ideal $x$ of $R_{\beta}$, then $s\in\overline{R}_{\beta}/\pi\overline{R}_{\beta}$. This is clear since the map $\overline{R}_{\beta}/\pi\overline{R}_{\beta}\to\overline{R}_{\beta}/\pi\overline{R}_{\beta}[f^{-1}]$ is injective by the regularity result proven above and using that $R_{\beta,x}\to\widehat{R}_{\beta,x}$ is faithfully flat. ∎
@@
Analogously define $\mathbf{Syn}^{c-\infty}(\mathcal{L},r)$ to be the object in the bounded derived category of sheaves on $X$ given by the total complex associated to the morphism of complexes $\iota^{c-\infty}\circ(1-\frac{\varphi}{p^{r}})\oplus\iota^{c-\infty}$ :
$\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\right)\to\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\right)\oplus\mathrm{R}\Gamma\left(X_{\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathrm{Gr}^{r}\mathbb{B}_{\mathrm{dR}}^{c-\infty}\right).$
@@
Let $\mathrm{H}^{i}\bigl(\mathbf{Syn}(\mathcal{L},r)\bigr)$, resp. $\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c-\infty}(\mathcal{L},r)\bigr)$ to be the $i$-th homology groups of the complexes $\mathbf{Syn}(\mathcal{L},r)$ and $\mathbf{Syn}^{c-\infty}(\mathcal{L},r)$ respectively.
@@
We have morphisms of complexes $\mathbf{Syn}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}(\mathcal{L},r)$ and $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c-\infty}(\mathcal{L},r)$.
@@
It follows from Theorem 1 and Definition 9 and by [DLL+23, Cor. 2.4.2] that have a quasi isomorphism
@@
compatible with Frobenius. The same holds with support conditions, see [LLZ23, Cor. 2.5.13] for the de Rham case and Theorem 1 in the crystalline case, noticing that the isomorphisms $\xi_{\mathrm{dR}}$ and $\xi_{\beta,\mathrm{crys}}$ of Definition 9 induce analogous and compatible isomorphisms $\xi_{\mathrm{dR}}^{c-\infty}$ and $\xi_{\beta,\mathrm{crys}}^{c-\infty}$ replacing $\mathcal{O}\mathbb{B}_{\mathrm{dR}}$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ with $\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c-\infty}$ and $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}$. It follows that we have morphisms of complexes